. Question Completion Status: Question 1 Safari Fusion reported shareholders’ equity as follows: Preferred stock, $10.00 par value Authorized 10,000 shares; issued–None $ — Common stock, $1 par value Authorized 30,000 shares; issued 15,000 shares 15,000 Capital in excess of par value 45,000 Retained earnings 90,000 Total paid-in capital 150,000 Less: Treasury stock, at cost (2,000 common shares) (16,000) Total stockholders’ equity $114,000 Requirements 1. What was the average issue price per share of the common stock? 2. Journalize the issuance of 1,000 shares of common stock at $4 per share. Use Safari’s account titles. 3. How many shares of Safari’s common stock are outstanding? 4. How many shares of common stock would be outstanding after Safari split its common stock 3 for 1? 5. Use Safari account titles, journalize the distribution of a 10% stock dividend when the market price of Safari common stock is $5 per share. Safari distributes the common stock dividend on the shares outstanding, which were computed in requirement 3. 6. Journalize the following treasury stock transactions, which occur in the order given: a. Safari purchases 500 shares of treasury stock at $8 per share. b. Safari sells 100 shares of treasury stock for $9 per share. c. Safari sells 200 shares of treasury stock for $6 per share. Path: p Words:0 10 points Click Save and Submit to save and submit. Click Save All Answers to save all answers.

   What was the average issue price per share of the common stock?

2.     Journalize the issuance of 1,000 shares of common stock at $4 per share. Use Safari’s account titles.

3.     How many shares of Safari’s common stock are outstanding?

4.     How many shares of common stock would be outstanding after Safari split its common stock 3 for 1?

5.     Use Safari account titles, journalize the distribution of a 10% stock dividend when the market price of Safari common stock is $5 per share. Safari distributes the common stock dividend on the shares outstanding, which were computed in requirement 3.

6.     Journalize the following treasury stock transactions, which occur in the order given:

a.   Safari purchases 500 shares of treasury stock at $8 per share.

b.   Safari sells 100 shares of treasury stock for $9 per share.

c.   Safari sells 200 shares of treasury stock for $6 per share.
Answer #1

1. Average Issue price of common stock = ($15000+45000)/15000 = $4 per share

2.

Account Titles Debit Credit
Cash $              4,000
       Common Stock $            1,000
       Paid in capital in excess of par $            3,000

3. Common Shares outstanding = 15000+1000-2000 = 14000 shares (If 1000 shares are issued afresh)
otherwise
Common Shares outstanding = 15000-2000 = 13000 shares (If 1000 shares are not issued extra)

4. Common stock after 3 for 1 split = 16000 x 3 = 48000 shares (If 1000 shares are issued afresh)
otherwise
Common stock after 3 for 1 split = 15000 x 3 = 45000 shares (If 1000 shares are not issued extra)

5. (If 1000 shares are issued afresh)

Account Titles Debit Credit
Retained Earnings $              7,000 =14000*10%*5
       Common Stock $            1,400 =14000*10%*1
       Paid in capital in excess of par $            5,600 =7000-1400

(If 1000 shares are not issued extra)

Account Titles Debit Credit
Retained Earnings $              6,500
       Common Stock $            1,300
       Paid in capital in excess of par $            5,200

6.

Account Titles Debit Credit
Treasury Stock $              4,000
        Cash $            4,000
Cash $                  900
     Treasury Stock $               800
     Paid in capital from treasury stock $               100
Cash $              1,200
Paid in capital from treasury stock $                  100
Retained Earnings $                  300
      Treasury Stock $            1,600